HETAV PATEL

Grade 11,

When a cieling fan is switched on it makes 10 revolutions in 1st 3 seconds Assumming a uniform angular acceleration how many rotations it will make in next 3 seconds

When a cieling fan is switched on it makes 10 revolutions in 1st 3 seconds Assumming a uniform angular acceleration how many rotations it will make in next 3 seconds

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear pallavi

θ= 1/2 αt2

10 =1/2 * α *9

α =20/9 rev /sec2

 

so revolution in 6 sec

θ= 1/2 αt2

  = 1/2 * 20/9 * 36

   =40

revolution in next 3 sec = revolution in 6 sec - revolution in 3 sec

                                  =40    -10

                                  =30

Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin